What is the surface area of a cylinder with a base diameter of 14 and a height of 16????!!!??
1
Expert's answer
2012-01-10T08:44:48-0500
surface area=base area+lateral surface area
base area=pi*radius^2=49*pi
lateral surface area=2*pi*height*radius=14*pi*16=224*pi
Hence
surface area=49*pi+224*pi=273*pi